# RemixFusion: Residual-based Mixed Representation for Large-scale Online RGB-D Reconstruction

### v1

@@ 3.3. Residual-based Bundle Adjustment
However, using the MLP to encode the complete camera poses only based on the frame index is inefficient and difficult for training. Inspired by the residual-based mixed representation (Section 3.2), we propose to change the output of from the complete pose to only the residual pose, aiming to reduce network learning complexity and thereby enhance training efficiency. Specifically, we encode the pose changes along with the frame index in . Note that the frame index should be normalized to according to the number of keyframes for better convergence. The residuals are added back to the initial poses to gain the final results . Compared to the independent optimization of each keyframe, the MLP is more comprehensive, and the residual-based BA is more globally aware and efficient. The process of the residual-based BA can be formulated as Eq. 16. The and correspond to and in Eq. 1 respectively. The residual-based BA is running with iterations every frames. denotes the normalization of the frame index for better convergence.
@@
Inspired by the Simulated Annealing algorithm (SA), gradient amplification (GA) seems like a good choice to overcome this problem. The key idea is to amplify the optimization gradients near the reconstructed surface and encourage the BA optimization to explore more. The simplest way to amplify the gradients is to move the cameras in the direction they are facing or in a random direction, by a certain distance. This would lead to misalignment of the currently observed surfaces and reconstructed surfaces. The lack of consideration of the 3D geometric structures causes the BA module to fail to achieve the expected results due to gradient amplification disorder.
@@
Figure 5(a) visualizes the optimization process of the bundle adjustment on the scene foobar of BS3D. The states, colorized by the errors of pose estimation, are 6D poses of all keyframes, which are visualized in 2D. The optimization starts with the poses of front-end tracking. The red and blue regions indicate high and low errors, respectively. BA with the proposed GA converges fast to the global minima, while BA without GA tends to be stuck in the local minima, and the improvement is marginal. The visualization is shown by the alignment of point clouds generated with the optimized poses and ground-truth poses, which correspond to the yellow and blue ones in Figure 5(b) . Overall, the proposed GA helps the bundle adjustment in large-scale scenes explore more and converge to the global minima. Extensive experiments show that the residual-based BA is efficient and robust, which can steadily refine the estimated initial camera poses with more global comprehension.
@@ 3.4. System Implementation
We illustrate the proposed scalable pose estimation for large-scale scenes using randomized optimization and system implementation, including sampling and keyframe selection, in this section.
@@
First, we review the core insight of randomized optimization used in (Zhang et al., 2021), as well as the basic principles and cost functions of pose estimation. In brief, the goal is to provide accurate 6DoF camera pose for live RGB-D frames . We first back-project the pixel with depth to its 3D position in the camera space, its point-to-surface distance in the world space could then be queried using . defines the difference between the queried TSDF values from the moving volume and the approximated ground-truth TSDF supervision, which is similar to Eq. 10 and Eq. 11. The minimization of these point-to-surface distances described in Eq. 18 is tackled by the randomized optimization with iterations using a pre-sampled particle swarm template (PST). We encourage readers to refer to supplementary materials for more details.
@@
One main difference of our method against ROSEFusion (Zhang et al., 2021) is that we leverage a moving volume to perform pose estimation to make our tracking module scalable to large-scale scenes. This is inspired by the success of the moving TSDF-Fusion strategy (Whelan et al., 2012; Roth and Vona, 2012). The process of the management of the moving explicit volume and pose estimation based on randomized optimization is illustrated in Algorithm 1. We initialize the first moving volume given the first camera position , and the input RGB-D frames , which are observed within volume are integrated. The camera pose would be marked as an anchor each time the volume is moved. We transform the moving volume only when the Euclidean distance between the current camera and the last anchor camera exceeds a threshold in any direction of the 3 axes. We create a new moving volume , inheriting the overlapping regions from to minimize the transferring costs, shown in Eq. 19. After swapping, is abandoned.
where refers to the voxel coordinate and is the default value of the TSDF volume. Here we set its TSDF value to one and the other attributes, including color and weights, to zero. In our implementation, we use a dense axis-aligned moving TSDF volume with resolution higher than that of . Our system can accurately perceive local environments with a fixed and moderate memory cost to maintain efficiency.
@@
Following Co-SLAM (Wang et al., 2023), we downsample each keyframe and only save 5 of the pixels in the keyframe database that contains the RGB, depth and pose data. The mapping process is continuously running every frame, and we sample pixels from the keyframe database each time. After sampling, we can back-project the pixels to 3D space, and sample points along the rays of different pixels by ray casting, which contains points uniformly sampled and sampled near the observed surface. Then, we can get the points , which are points in total to be queried and learned.
@@ 4.1. Experimental Setup
We evaluate our method on two large-scale indoor datasets BS3D (Mustaniemi et al., 2023) and uHumans2 (Rosinol et al., 2020). We also provide the comparison for the room-level publicly available dataset: Replica (Straub et al., 2019), TUM RGB-D (Sturm et al., 2012), ScanNet (Dai et al., 2017a), and FastCaMo (Zhang et al., 2021). This evaluation aims to benchmark our tracking and reconstruction performance against current state-of-the-art methods. BS3D is a challenging real-world large-scale RGB-D dataset, comprising over 10,000 RGB-D images annotated with ground-truth trajectories from a motion capture system. uHumans2 is a large-scale publicly available synthetic dataset with multiple rooms and complicated layouts, presenting a significant challenge for SLAM. We evaluate eight large-scale sequences of BS3D, two sequences of uHumans2, alongside commonly referenced sequences from Replica, ScanNet, TUM RGB-D, and FastCaMo-Synth (noise-free). Furthermore, additional qualitative comparisons are conducted on BS3D, uHumans2, FastCaMo-Large proposed in MIPS-Fusion (Tang et al., 2023), and self-captured sequences to highlight the different performance on large-scale scenes.
@@
In terms of 3D reconstruction quality, we follow (Zhu et al., 2022) and adopt Accuracy(cm), Completion(cm), and Completion ratio(%) with the threshold of 5cm and 10cm as well as F1-score. Following (Azinović et al., 2022; Zhu et al., 2022), we filter the noisy points that are not visible in any observation with ground-truth depth images for a fair comparison, as many neural approaches tend to predict numerous noisy points in empty spaces. Then we use the Iterative Closest Point (ICP) for alignment with ground-truth meshes. Metrics like PSNR, SSIM, LPIPS, and Depth-L1 (m) for 2D rendering comparison are adopted following (Keetha et al., 2024). As for camera tracking, the ATE RMSE (Sturm et al., 2012) is adopted. The system FPS (frames per second) and GPU memory usage are considered for efficiency comparison.
@@
Baselines
We compare our method to both the implicit and explicit state-of-the-art methods. The former includes NeRF-based methods like iMAP (Sucar et al., 2021), NICE-SLAM (Zhu et al., 2022), Co-SLAM (Wang et al., 2023), ESLAM (Johari et al., 2023) and MIPS-Fusion (Tang et al., 2023). The latter indicates ElasticFusion (Whelan et al., 2015), BundleFusion (Dai et al., 2017c) and BAD-SLAM (Schops et al., 2019). RTG-SLAM (Peng et al., 2024), GS-ICP SLAM (Ha et al., 2024), SplaTAM (Keetha et al., 2024), MonoGS (Matsuki et al., 2024), Photo-SLAM (Huang et al., 2024a) and LoopSplat (Zhu et al., 2025) are encompassed as the representative of the SOTA 3DGS-based SLAM approaches. For more comprehensive comparisons, the sparse SLAM framework ORB-SLAM3 (Campos et al., 2021) and DROID-SLAM (Teed and Deng, 2021) that specialize in robust camera tracking are included. The implementation of iMAP* is from NICE-SLAM. ESLAM, SplaTAM, and MonoGS are evaluated using images at half resolution as inputs on BS3D and uHumans2 datasets due to excessive GPU occupancy. For RTG-SLAM, a modified version was employed due to its high GPU memory allocation in global optimization on these two datasets. More details can be found in the supplementary materials.
@@ 4.2. Quantitative and Qualitative Comparison
In this section, we present quantitative and qualitative results, including the camera pose estimation, 3D mesh reconstruction, and 2D rendering on different datasets. The lightweight version of our method, requiring fewer optimization iterations for both tracking and mapping, is denoted as RemixFusion-lite. Details can be found in the supplementary materials.
@@
Table 2 presents the results of 3D reconstruction using ground-truth poses to eliminate the tracking effects in SLAM. Thanks to the residual-based mixed representation, our method is the most competitive one in accuracy (4.34cm) and completeness (3.56cm), surpassing the state-of-the-art by 5 in accuracy and 9 in completeness. The second-best method, after our own, is Co-SLAM, which achieves the accuracy of 4.57cm and the completeness of 3.91cm. While the completeness ratio of ESLAM is the second-best for implicit methods, its completeness and accuracy remain suboptimal, indicating the incompact reconstruction. Notably, with ground-truth poses, the best 3DGS-based method demonstrates inferior reconstruction performance (31.9 worse than the best implicit method in accuracy) with the online setting. RTG-SLAM leverages an efficient and compact 3DGS representation to save memory, but the reconstruction performance is not desirable, indicating the trade-off between efficiency and accuracy. Although GS-ICP SLAM has a comparable mapping FPS to our method, its reconstruction accuracy is 4.75 worse than ours. This indicates that achieving high-quality geometric reconstruction with 3DGS requires significantly more optimization iterations in real-world large-scale scenarios, which may lead to compromised 3D reconstruction under real-time SLAM constraints. Note that the outputs of MIPS-Fusion are obtained by running the official code, and some objects are missing. This leads to worse performance of completeness and comparable performance of completeness ratio. The lightweight version of our method only performs mapping with a few optimization iterations in this experiment and is significantly faster, with only a slight decrease in reconstruction accuracy (-0.36cm) and completeness (-0.5cm), which further proves the effectiveness and robustness of our method.
@@
Table 1 shows ATE RMSE (cm) on 8 sequences of BS3D for our method compared to five state-of-the-art implicit RGB-D SLAM and three explicit traditional methods. Evaluations of the state-of-the-art 3DGS-based methods (SplaTAM, MonoGS, RTG-SLAM, GS-ICP SLAM, LoopSplat and Photo-SLAM) and the sparse SLAM (ORB-SLAM3) are also incorporated for a more comprehensive comparison for tracking accuracy.
Our method attains the best tracking performance (4.6cm) for 8 challenging sequences on average while running in real time. There is 28.1 improvement for our method in tracking on average compared to the SOTA methods. Implicit methods struggle with pose estimation in these large-scale scenes. Specifically, implicit methods, except for NICE-SLAM and ESLAM , all fail in tracking on the scene cafeteria, which is more than 400 and the layouts are complex, even including some textureless areas. These factors pose challenges for robust camera pose estimation and real-time reconstruction, resulting in unsatisfactory reconstruction quality, unstable pose estimation, and inadequate bundle adjustment for most methods. Although our method is inferior to the sparse SLAM ORB-SLAM3 on this challenging sequence, our method surpassed ESLAM by 11.7. For the scene hub, our method underperforms the best method by a margin of 0.2cm. Although NICE-SLAM and ESLAM demonstrate stable tracking performance on BS3D, they suffer from low FPS and high GPU memory consumption in large-scale scenes, as shown in Table 8. In contrast, our method maintains accuracy and robustness on these large-scale scenes RemixFusion-lite (over 25FPS), the lightweight version, is the second-best alternative on average, further proving the efficiency and robustness of our method.
@@
NICE-SLAM | 14.49 | 21.8 | 0.929 | 0.395 | 0.134 | 6.79 | 25.35 | 0.976 | 0.345 | 0.086 | 10.64 | 23.575 | 0.953 | 0.370 | 0.110 |
Co-SLAM | 1268.23 | 17.79 | 0.838 | 0.752 | 1.885 | 26.08 | 0.977 | 0.412 | 0.171 | 10.95 | 639.59 | 21.94 | 0.908 | 0.582 | 1.028 |
@@
Explicit dense RGB-D methods struggle with robustness in these challenging scenes. Among them, only ElasticFusion, RTG-SLAM GS-ICP SLAM, Photo-SLAM, and ORB-SLAM3 consistently succeed in tracking on all scenes, while BundleFusion and BAD-SLAM encounter failures in certain cases, primarily due to front-end odometry issues, which is also validated in (Mustaniemi et al., 2023). Although ElasticFusion can finish tracking on these sequences, its trajectory proves inaccurate.
3DGS-based methods, including RTG-SLAM, GS-ICP SLAM, and Photo-SLAM, are comparable to NICE-SLAM in tracking. These methods leverage either multi-level ICP or ORB features (Mur-Artal and Tardós, 2017) for tracking, which is decoupled from 3DGS optimization. In contrast, other 3DGS-based methods, including SplaTAM, MonoGS, and LoopSplat, which leverage the rendering loss as the key objective function for camera pose optimization, exhibit unstable tracking performance on this large-scale dataset. Although they can attain high-fidelity rendering given good camera poses, it is challenging to optimize the 3DGS and camera poses simultaneously in large-scale scenes, since there are drastically more 3DGS to optimize. Notably, LoopSplat achieves successful tracking on the foobar and lounge sequences but fails on the remaining sequences on BS3D. SplaTAM fails on 3 sequences and MonoGS fails on 1 sequence. In contrast, our method is robust and surpasses the best 3DGS-based method (RTG-SLAM) by 62.3 for tracking.
@@
iMAP∗ | 70.1 | 4.5 | 2.2 | 2.3 | 1.7 | 0.5 | 58.4 | 2.6 | 17.8 | 0.2 | 56.0 | 32.1 | 17.5 | 70.5 | 32.1 | 11.9 | 36.7 | 0.2 | 7.2 | 2.1 | 9.0 | 6.1 | 0.1 | 20.2 |
NICE-SLAM | 1.7 | 2.0 | 1.6 | 1.0 | 0.9 | 1.4 | 4.0 | 3.1 | 2.0 | 1.1 | 8.6 | 12.3 | 8.1 | 10.3 | 12.9 | 5.6 | 9.6 | 0.6 | 2.7 | 1.8 | 3.0 | 2.5 | 0.2 | 4.7 |
@@
ESLAM | 0.7 | 0.7 | 0.5 | 0.6 | 0.6 | 0.6 | 0.7 | 0.6 | 0.6 | 7.3 | 7.3 | 8.5 | 7.5 | 6.5 | 9.0 | 5.7 | 7.4 | 2.3 | 2.5 | 1.1 | 2.4 | 2.0 | 0.2 | 3.4 |
MIPS-Fusion | 1.1 | 1.2 | 1.1 | 0.7 | 0.8 | 1.3 | 2.2 | 1.1 | 1.2 | 2.8 | 7.9 | 10.7 | 9.7 | 9.7 | 14.2 | 7.8 | 10.0 | 3.1 | 3.0 | 1.4 | 4.6 | 3.0 | 3.0 | 4.7 |
@@
MonoGS | 0.3 | 0.2 | 0.3 | 0.4 | 0.2 | 0.3 | 0.1 | 0.8 | 0.3 | 0.7 | 9.8 | 32.1 | 8.9 | 10.7 | 21.8 | 7.9 | 15.2 | 1.7 | 1.6 | 1.4 | 1.5 | 1.5 | 1.4 | 5.7 |
SplaTAM | 0.3 | 0.4 | 0.3 | 0.5 | 0.3 | 0.3 | 0.3 | 0.6 | 0.4 | 0.3 | 12.8 | 10.1 | 17.7 | 12.1 | 11.1 | 7.5 | 11.9 | 0.5 | 3.4 | 1.2 | 5.2 | 3.3 | 0.3 | 5.2 |
LoopSplat | 0.3 | 0.2 | 0.2 | 0.2 | 0.2 | 0.5 | 0.2 | 0.3 | 0.3 | 0.5 | 6.2 | 7.1 | 7.4 | 10.6 | 8.5 | 6.6 | 7.7 | 0.6 | 2.1 | 1.6 | 3.2 | 2.3 | 0.6 | 3.4 |
Photo-SLAM | 0.5 | 0.4 | 0.3 | 0.5 | 0.4 | 1.3 | 0.8 | 0.6 | 0.6 | 28.6 | 8.3 | 6.7 | 9.2 | 8.5 | 78.8 | 7.5 | 19.8 | 22.3 | 2.6 | 0.3 | 1.0 | 1.3 | 21.3 | 7.3 |
RTG-SLAM | 0.2 | 0.2 | 0.1 | 0.2 | 0.1 | 0.2 | 0.2 | 0.2 | 0.2 | 5.9 | 125.1 | 109.1 | 128.8 | 7.8 | 28.1 | 6.9 | 67.6 | 2.9 | 1.7 | 0.4 | 1.1 | 1.1 | 3.8 | 23.0 |
GS-ICP SLAM | 0.2 | 0.2 | 0.1 | 0.2 | 0.1 | 0.2 | 0.2 | 0.2 | 0.2 | 22.4 | 78.3 | 94.5 | 41.2 | 112.6 | 59.8 | 20.1 | 67.8 | 24.3 | 2.7 | 1.8 | 2.7 | 2.4 | 24.2 | 23.5 |
RemixFusion | 0.5 | 0.4 | 0.3 | 0.2 | 0.5 | 0.5 | 0.4 | 0.4 | 0.4 | 14.3 | 6.9 | 10.3 | 9.3 | 6.7 | 15.4 | 7.6 | 9.4 | 12.0 | 2.3 | 1.8 | 2.4 | 2.2 | 10.7 | 4.0 |

### v3

@@ 3.3. Residual-based Bundle Adjustment
However, using the MLP to encode the complete camera poses only based on the frame index is inefficient and difficult for training (Eq. 15). Inspired by the residual-based mixed representation (Section 3.2) and (Bian et al., 2024), we propose to change the output of from the complete pose to only the residual pose, aiming to reduce network learning complexity and thereby enhance training efficiency. Specifically, we encode the pose changes along with the frame index in . Note that the frame index should be normalized to according to the number of keyframes for better convergence. The residuals are added back to the initial poses to gain the final results . Compared to the independent optimization of each keyframe, the MLP is more comprehensive, and the residual-based BA is more globally aware and efficient. The residual-based BA can be formulated as Eq. 16. The and correspond to and in Eq. 1 respectively. The residual-based BA is running with iterations every frames. denotes the normalization of the frame index.
@@
Inspired by the Simulated Annealing algorithm (SA), gradient amplification (GA) seems like a good choice to overcome this problem. The key idea is to amplify the optimization gradients near the reconstructed surface and encourage the BA optimization to explore more. The simplest way to amplify the gradients is to move the cameras in the direction they are facing or in a random direction by a certain distance. This would lead to misalignment of the currently observed surfaces and reconstructed surfaces. The lack of consideration of the 3D geometric structures makes the BA module hard to achieve the expected results due to gradient amplification disorder.
@@
Figure 5(a) visualizes the optimization process of the bundle adjustment on the scene foobar of BS3D. The states, colorized by the errors of pose estimation, are 6D poses of all keyframes, which are visualized in 2D. The optimization starts with the poses of front-end tracking. Red denotes higher errors, while blue denotes lower errors. BA with the proposed GA converges fast to the global minima, whereas BA without GA tends to be stuck in local minima and yields only marginal improvement. The visualization is shown by the alignment of point clouds generated with the optimized poses and ground-truth poses, which correspond to the yellow and blue ones in Figure 5(b). Overall, the proposed GA helps the bundle adjustment in large-scale scenes explore more and converge to the global minima. Extensive experiments show that the residual-based BA is efficient and robust, which can steadily refine the initially estimated camera poses with more global comprehension.
@@ 3.4. System Implementation
In this section, we illustrate the proposed scalable pose estimation for large-scale scenes using randomized optimization and system implementation, including sampling and keyframe selection.
@@
First, we review the core insight of randomized optimization used in (Zhang et al., 2021), as well as the basic principles and cost functions of pose estimation. In brief, the goal is to provide accurate 6DoF camera poses for live RGB-D frames . We first back-project the pixel with depth to its 3D position in the camera space, its point-to-surface distance in the world space could then be queried using and the transformation . defines the difference between the queried TSDF values from the moving volume and the approximated ground-truth TSDF supervision, which is similar to Eq. 10 and Eq. 11. The minimization of these point-to-surface distances described in Eq. 18 is tackled by the randomized optimization with iterations using a pre-sampled particle swarm template (PST). We encourage readers to refer to supplementary materials for more details.
@@
One main difference of our method against ROSEFusion (Zhang et al., 2021) is that we leverage a moving volume to perform pose estimation to make our tracking module scalable to large-scale scenes. This is inspired by the success of the moving TSDF-Fusion strategy (Whelan et al., 2012; Roth and Vona, 2012). The process of the management of the moving explicit volume and pose estimation based on randomized optimization is illustrated in Algorithm 1. Given the first camera position and the input RGB-D frames , we initialize the first moving volume . Only the parts observed within the volume are integrated. The camera pose would be marked as an anchor each time the volume is moved. We transform the moving volume only when the Euclidean distance between the current camera and the last anchor camera exceeds a threshold in any direction of the 3 axes. We create a new moving volume , inheriting the overlapping regions from to minimize the transferring costs, shown in Eq. 19.
where refers to the world coordinate and is the default value of the TSDF volume. Here, we set its TSDF value to one and the other attributes, including color and weights, to zero.
After swapping, is marked as the new anchor volume . In our implementation, we use a dense axis-aligned moving TSDF volume with resolution higher than that of . Our system can accurately perceive local environments with a fixed and moderate memory cost to maintain efficiency.
@@
Following Co-SLAM (Wang et al., 2023), we downsample each keyframe and only save 5 of the pixels in the keyframe database that contains the RGB, depth, and pose data. The mapping process is continuously running every frame, and we sample pixels from the keyframe database each time. After sampling, we can back-project the pixels to 3D space, and sample points along the rays of different pixels by ray casting, including uniformly sampled points and points sampled near the observed surface. Then, we can get the points , which are points in total to be queried and learned.
@@ 4.1. Experimental Setup
We evaluate methods on two large-scale indoor datasets BS3D (Mustaniemi et al., 2023) and uHumans2 (Rosinol et al., 2020). We also provide the comparison for the room-level publicly available dataset: Replica (Straub et al., 2019), TUM RGB-D (Sturm et al., 2012), ScanNet (Dai et al., 2017a), and FastCaMo (Zhang et al., 2021). This evaluation aims to benchmark our tracking and reconstruction performance against current state-of-the-art methods. BS3D is a challenging real-world large-scale RGB-D dataset, comprising over 10,000 RGB-D images annotated with ground-truth trajectories from a motion capture system. uHumans2 is a large-scale publicly available synthetic dataset with multiple rooms and complicated layouts, presenting a significant challenge for SLAM. We evaluate eight large-scale sequences of BS3D, two sequences of uHumans2, alongside commonly referenced sequences from Replica, ScanNet, TUM RGB-D, and FastCaMo-Synth (noise-free). Furthermore, additional qualitative comparisons are conducted on BS3D, uHumans2, FastCaMo-Large proposed in MIPS-Fusion (Tang et al., 2023), and self-captured sequences to highlight the different performance on large-scale scenes.
@@
In terms of 3D reconstruction quality, we follow (Zhu et al., 2022) and adopt Accuracy(cm), Completion(cm), and Completion ratio(%) at the threshold of 5cm and 10cm, as well as the F1-score. Following (Azinović et al., 2022; Zhu et al., 2022), we filter the noisy points that are not visible in any observation with ground-truth depth images for a fair comparison, as many neural approaches tend to predict numerous noisy points in empty spaces. Then we use the Iterative Closest Point (ICP) for alignment with ground-truth meshes. Metrics like PSNR, SSIM, LPIPS, and Depth-L1 (m) for 2D rendering comparison are adopted following (Keetha et al., 2024). As for camera tracking, the ATE RMSE (Sturm et al., 2012) is adopted. The system FPS (frames per second) and GPU memory usage are considered for efficiency comparison.
Baselines
We compare our method to both the implicit and explicit state-of-the-art methods. The former includes NeRF-based methods like iMAP (Sucar et al., 2021), NICE-SLAM (Zhu et al., 2022), Co-SLAM (Wang et al., 2023), ESLAM (Johari et al., 2023) and MIPS-Fusion (Tang et al., 2023). The latter indicates ElasticFusion (Whelan et al., 2015), BundleFusion (Dai et al., 2017c) and BAD-SLAM (Schops et al., 2019). RTG-SLAM (Peng et al., 2024), GS-ICP SLAM (Ha et al., 2024), SplaTAM (Keetha et al., 2024), MonoGS (Matsuki et al., 2024), Photo-SLAM (Huang et al., 2024a) and LoopSplat (Zhu et al., 2025) are encompassed as the representative of the SOTA 3DGS-based SLAM approaches. For more comprehensive comparisons, the sparse SLAM ORB-SLAM3 (Campos et al., 2021) and DROID-SLAM (Teed and Deng, 2021) that specializes in robust camera tracking are included. The implementation of iMAP* is from NICE-SLAM. ESLAM, SplaTAM, and MonoGS are evaluated using images at half resolution as inputs on BS3D and uHumans2 datasets due to out-of-memory issues. For RTG-SLAM, a modified version was employed due to its high GPU memory allocation in global optimization on these two datasets. More details can be found in the supplementary materials.
@@ 4.2. Quantitative and Qualitative Comparison
In this section, we present quantitative and qualitative results, including the camera pose estimation, 3D mesh reconstruction, and 2D rendering on different datasets. The lightweight version of our method, requiring fewer optimization iterations, is denoted as RemixFusion-lite. Details can be found in the supplementary materials.
@@
Table 2 is the result of 3D reconstruction using ground-truth poses to eliminate the tracking effects in SLAM. Thanks to the residual-based mixed representation, our method is the most competitive one in accuracy (4.34cm) and completeness (3.56cm), surpassing the state-of-the-art by 5 in accuracy and 9 in completeness. The second-best method, after our own, is Co-SLAM, which achieves the accuracy of 4.57cm and the completeness of 3.91cm. While the completeness ratio of ESLAM is the second-best for implicit methods, its completeness and accuracy remain suboptimal, indicating the incomplete reconstruction of its tri-plane representation.
Notably, with ground-truth poses, the best 3DGS-based method demonstrates inferior reconstruction performance (46.8 worse than the best implicit method in accuracy) with the online setting. RTG-SLAM leverages an efficient and compact 3DGS representation to save memory, but the reconstruction performance is not desirable, indicating the trade-off between efficiency and accuracy. Although GS-ICP SLAM has a comparable mapping FPS to our method, its reconstruction accuracy is 4.75cm worse than ours. This indicates that achieving high-quality geometric reconstruction with 3DGS requires significantly more optimization iterations in real-world large-scale scenarios, which may lead to compromised 3D reconstruction under real-time SLAM constraints. Note that the outputs of MIPS-Fusion are obtained by running the official code, and some objects are missing. This leads to worse performance of completeness and comparable performance of completeness ratio. The lightweight version of our method only performs mapping with a few optimization iterations in this experiment and is significantly faster, with only a slight decrease in reconstruction accuracy (-0.36cm) and completeness (-0.5cm), which further proves the effectiveness and robustness of our method.
@@
Table 1 shows ATE RMSE (cm) on 8 sequences of BS3D for our method compared to five state-of-the-art implicit methods and three explicit traditional methods. Evaluations of the state-of-the-art 3DGS-based methods (SplaTAM, MonoGS, RTG-SLAM, GS-ICP SLAM, LoopSplat and Photo-SLAM) and the sparse SLAM (ORB-SLAM3) are also incorporated for a more comprehensive comparison for tracking accuracy.
Our method attains the best tracking performance (4.6cm) for 8 challenging sequences on average while running in real time. There is 28.1 improvement for our method in tracking on average compared to the SOTA methods. Implicit methods struggle with pose estimation in these large-scale scenes. Specifically, implicit methods, except for NICE-SLAM and ESLAM, all struggle in tracking on the scene cafeteria, which is more than 400. The layouts are complex, even including some textureless areas. These factors pose challenges for robust camera pose estimation and real-time reconstruction, resulting in unsatisfactory reconstruction quality, unstable pose estimation, and inadequate bundle adjustment for most methods. Although our method is inferior to the sparse SLAM ORB-SLAM3 on this challenging sequence, our method surpasses ESLAM by 11.7. For the scene hub, our method indicates a marginal performance gap of 0.2cm between our method and the best method. Although NICE-SLAM and ESLAM demonstrate stable tracking performance on BS3D, they suffer from low FPS and high GPU memory consumption in large-scale scenes, as shown in Table 8. In contrast, our method maintains accuracy and robustness on these large-scale scenes. RemixFusion-lite (over 25FPS), the lightweight version, is the second-best alternative on average, further proving the efficiency and robustness of our method.
@@
| NICE-SLAM | 14.49 | 21.80 | 0.929 | 0.395 | 0.134 | 6.79 | 25.35 | 0.976 | 0.345 | 0.086 | 10.64 | 23.58 | 0.953 | 0.370 | 0.110 |
| Co-SLAM | 1268.23 | 17.79 | 0.838 | 0.752 | 1.885 | 10.95 | 26.08 | 0.977 | 0.412 | 0.171 | 639.59 | 21.94 | 0.908 | 0.582 | 1.028 |
@@
Explicit dense RGB-D methods struggle with robustness in these challenging scenes. Among them, only ElasticFusion, RTG-SLAM, GS-ICP SLAM, Photo-SLAM, and ORB-SLAM3 consistently succeed in tracking on all scenes. While BundleFusion and BAD-SLAM encounter failures in certain cases, primarily due to front-end odometry issues, which are also validated in (Mustaniemi et al., 2023). Although ElasticFusion can finish tracking on these sequences, its trajectory proves inaccurate.
3DGS-based methods, including RTG-SLAM, GS-ICP SLAM, and Photo-SLAM, are comparable to NICE-SLAM in tracking. These methods leverage either multi-level ICP or ORB features (Mur-Artal and Tardós, 2017) for tracking, which is decoupled from 3DGS optimization. In contrast, other 3DGS-based methods, including SplaTAM, MonoGS, and LoopSplat, which leverage the rendering loss as the key objective function for camera pose optimization, exhibit unstable tracking performance on this large-scale dataset. Although they can attain high-fidelity rendering given good camera poses, it is challenging to optimize the 3DGS and camera poses simultaneously in large-scale scenes, since there are drastically more 3DGS to optimize. Notably, LoopSplat achieves successful tracking on the foobar and lounge sequences but fails on the remaining sequences on BS3D. SplaTAM fails on 3 sequences, and MonoGS fails on 1 sequence. In contrast, our method is robust and surpasses the best 3DGS-based method (RTG-SLAM) by 62.3 for tracking.
@@
| iMAP∗ | 70.1 | 4.5 | 2.2 | 2.3 | 1.7 | 0.5 | 58.4 | 2.6 | 17.8 | 0.2 | 56.0 | 32.1 | 17.5 | 70.5 | 32.1 | 11.9 | 36.7 | 0.2 | 7.2 | 2.1 | 9.0 | 6.1 | 0.1 | 22.4 |
| NICE-SLAM | 1.7 | 2.0 | 1.6 | 1.0 | 0.9 | 1.4 | 4.0 | 3.1 | 2.0 | 1.1 | 8.6 | 12.3 | 8.1 | 10.3 | 12.9 | 5.6 | 9.6 | 0.6 | 2.7 | 1.8 | 3.0 | 2.5 | 0.2 | 4.8 |
@@
| ESLAM | 0.7 | 0.7 | 0.5 | 0.6 | 0.6 | 0.6 | 0.7 | 0.6 | 0.6 | 7.3 | 7.3 | 8.5 | 7.5 | 6.5 | 9.0 | 5.7 | 7.4 | 2.3 | 2.5 | 1.1 | 2.4 | 2.0 | 0.2 | 3.3 |
| MIPS-Fusion | 1.1 | 1.2 | 1.1 | 0.7 | 0.8 | 1.3 | 2.2 | 1.1 | 1.2 | 2.8 | 7.9 | 10.7 | 9.7 | 9.7 | 14.2 | 7.8 | 10.0 | 3.1 | 3.0 | 1.4 | 4.6 | 3.0 | 3.0 | 4.6 |
@@
| MonoGS | 0.3 | 0.2 | 0.3 | 0.4 | 0.2 | 0.3 | 0.1 | 0.8 | 0.3 | 0.7 | 9.8 | 32.1 | 8.9 | 10.7 | 21.8 | 7.9 | 15.2 | 1.7 | 1.6 | 1.4 | 1.5 | 1.5 | 1.4 | 5.8 |
| SplaTAM | 0.3 | 0.4 | 0.3 | 0.5 | 0.3 | 0.3 | 0.3 | 0.6 | 0.4 | 0.3 | 12.8 | 10.1 | 17.7 | 12.1 | 11.1 | 7.5 | 11.9 | 0.5 | 3.4 | 1.2 | 5.2 | 3.3 | 0.3 | 4.9 |
| LoopSplat | 0.3 | 0.2 | 0.2 | 0.2 | 0.2 | 0.5 | 0.2 | 0.3 | 0.3 | 0.5 | 6.2 | 7.1 | 7.4 | 10.6 | 8.5 | 6.6 | 7.7 | 0.6 | 2.1 | 1.6 | 3.2 | 2.3 | 0.6 | 3.3 |
| Photo-SLAM | 0.5 | 0.4 | 0.3 | 0.5 | 0.4 | 1.3 | 0.8 | 0.6 | 0.6 | 28.6 | 8.3 | 6.7 | 9.2 | 8.5 | 78.8 | 7.5 | 19.8 | 22.3 | 2.6 | 0.3 | 1.0 | 1.3 | 21.3 | 7.5 |
| RTG-SLAM | 0.2 | 0.2 | 0.1 | 0.2 | 0.1 | 0.2 | 0.2 | 0.2 | 0.2 | 5.9 | 125.1 | 109.1 | 128.8 | 7.8 | 28.1 | 6.9 | 67.6 | 2.9 | 1.7 | 0.4 | 1.1 | 1.1 | 3.8 | 24.1 |
| GS-ICP SLAM | 0.2 | 0.2 | 0.1 | 0.2 | 0.1 | 0.2 | 0.2 | 0.2 | 0.2 | 22.4 | 78.3 | 94.5 | 41.2 | 112.6 | 59.8 | 20.1 | 67.8 | 24.3 | 2.7 | 1.8 | 2.7 | 2.4 | 24.2 | 24.4 |
| RemixFusion | 0.5 | 0.4 | 0.3 | 0.2 | 0.5 | 0.5 | 0.4 | 0.4 | 0.4 | 14.3 | 6.9 | 10.3 | 9.3 | 6.7 | 15.4 | 7.6 | 9.4 | 12.0 | 2.3 | 1.8 | 2.4 | 2.2 | 10.6 | 3.9 |